# A Superconducting Single-Atom Phonon Laser

### v1

@@ A Superconducting Single-Atom Phonon Laser
${}^{1}$avli Institute of Nanoscience, Delft University of Technology, PO Box 5046, 2600 GA Delft, The Netherlands
${}^{2}$ (UMR 7006), Université de Strasbourg, 67000 Strasbourg, France
${}^{3}$nstitute for Theory of Condensed Matter, Karlsruhe Institute of Technology, 76131 Karlsruhe, Germany
${}^{4}$nstitute for Quantum Materials and Technology, Karlsruhe Institute of Technology, 76344 Eggenstein-Leopoldshafen, Germany
${}^{\dagger}$mail: c.a.potts@tudelft.nl
$*{}^{*}$mail: g.a.steele@tudelft.nl
The development of quantum acoustics has enabled the cooling of mechanical objects to their quantum ground state, generation of mechanical Fock-states, and Schrödinger cat states. Such demonstrations have made mechanical resonators attractive candidates for quantum information processing, metrology, and tests of quantum gravity theories. Here, we experimentally demonstrate a direct quantum-acoustic equivalent of a single-atom laser. A single superconducting qubit coupled to a high-overtone bulk acoustic resonator is used to drive the onset of phonon lasing. We observe the absence of a sharp lower lasing threshold and characteristic upper lasing threshold, unique predictions of single-atom lasing. Lasing of an object with an unprecedented $25$ $\mu$g mass represents a new regime of laser physics and provides a foundation for integrating phonon lasers with on-chip devices.
@@ INTRODUCTION
The development of the laser in the early 1960s has proven to be one of the most important breakthroughs in modern physics. The laser found its niche in the advancement of atomic physics, resulting in numerous groundbreaking discoveries such as spectroscopic measurements of atomic species Schawlow (1982, 1978), laser cooling of atoms Phillips and Metcalf (1982); Phillips (1998), and the realization of Bose-Einstein condensation Anderson et al. (1995); Davis et al. (1995). Beyond this, lasers are indispensable tools in areas like gravitational wave detection Abbott et al. (2016), biology Chaurasiya et al. (2016), chemistry McCreery (2005), and even in medicine Khalkhal et al. (2019).
The advancements enabled by optical lasers have motivated the development of lasing states of mechanical vibrations — phonon lasersMahboob et al. (2013); Vahala et al. (2009); Pettit et al. (2019); Kuang et al. (2023); Behrle et al. (2023). Coherent sources of phonons are interesting for coupling mechanical strain to electron spins and color centers Shandilya et al. (2021), coherent driving of magnetic spin-waves Hatanaka et al. (2022, 2023), improving the resolution of force sensors Cui et al. (2021), fundamental tests of quantum mechanics Behrle et al. (2023), and nonlinear phononics Kuang et al. (2023). However, to date, phonon lasing has been realized with the center of mass motion of levitated microspheres or ions Vahala et al. (2009); Pettit et al. (2019); Kuang et al. (2023); Behrle et al. (2023), limiting the integrability of these coherent phonon sources. In other examples of phonon lasers in literature, the lasing state is the result of dispersive or dissipative optomechanical coupling Shandilya et al. (2021); Cui et al. (2021); Zhang et al. (2018); Grudinin et al. (2010); Potts et al. (2021); Parsa et al. (2023) — without population inversion — which are well described as mechanical parametric instabilities Kippenberg and Vahala (2008) or as optically driven pulsed semiconductor heterostructures Bron and Grill (1978); Wallentowitz et al. (1996); Camps et al. (2001).
@@
Figure 1: Schematic of the on-chip $\hbar$BAR device. (A) Rendering of the $\hbar$BAR device. The $\hbar$BAR device comprises two chips bounded in a flip-chip orientation. The top chip is 650 $\mu$m of sapphire, hosts the HBAR modes (red) and is coupled to the superconducting antenna using an aluminum nitride pad (red). The pocket-style transmon qubit (silver) is fabricated from NbTiN on the bottom silicon chip and coupled to the feedline via an on-chip readout resonator (blue).
@@
Figure 2: Qubit-induced acoustic linewidth narrowing. (A) Definition of modes and the coupling rates between the modes. The readout resonator $\hat{a}$ has input-output modes labelled $1$ and $2$. (B) Measured two-tone spectroscopy for drive power of -12.0 dBm set at room temperature. (C) Measured phonon-induced transparency window as a function of qubit drive power. Starting from the lowest curve, drive powers are -12.0, -6.0, and -2.0 dBm set at room temperature, respectively. With increasing power, two features can be noticed. The transparency window reduces in depth, and the full-width half-maximum linewidth narrows. (D) Extracted experimental full-width half-maximum of the transparency window as a function of qubit drive power. Colored points match the corresponding curves in (C).
@@
This article demonstrates, for the first time, the direct phononic analog of a single-atom optical laser McKeever et al. (2003); Dubin et al. (2010); An et al. (1994). A single two-level artificial atom driven coherently into population inversion couples to a bulk on-chip phonon mode generating a lasing state. The lasing is observed through a dramatically extended ringdown of the artificial atom excited by the phonon laser and demonstrates the nearly threshold-free nature of single-atom lasers. We also observe a unique predicted feature of single-atom lasers, an upper lasing threshold Dubin et al. (2010); Mu and Savage (1992); Ashhab et al. (2009). Our experimental observations agree with the theoretical and numerical models we use to characterize the system. Moreover, the bulk nature of the phonon mode may allow the future integration of color centers or quantum dots, enabling strain engineering or coherent mechanical driving. The coherent state generated via phonon lasing may be used as an efficient displacement pulse for generating large displacement Schrodinger cat states Bild et al. (2023). Generating larger cat states may allow for future tests of exotic theories of gravitationally induced quantum decoherence Gely and Steele (2021); Schrinski et al. (2023). Future implementations of this architecture may even enable phonon spectroscopy of vibrational transitions of complex molecules and crystal structures Wybourne and Wigmore (1988) or even biological samples Romero-Isart et al. (2011).
@@
Our device comprises a flip-chip $\hbar$BAR architecture with two bonded device chips Bild et al. (2023); von Lüpke et al. (2023, 2022); Chu et al. (2018, 2017). We have implemented a fully on-chip integration, which can multiplex different devices on a single silicon chip Välimaa et al. (2022); Crump, Välimaa, and Sillanpää (2023). A single feedline is coupled to the individual superconducting fixed-frequency transmon qubits. The transmon qubits were fabricated from niobium-titanium nitride for the bulk structures with aluminum Josephson junctions. A sapphire substrate, $650$ $\mu$m thick, was positioned above each transmon qubit and bonded to the silicon substrate, see Fig. 1. The sapphire chip supports a set of longitudinal high-overtone bulk acoustic wave resonances (HBARs) separated by a free spectral range of $8.54$ MHz. The electric field of a transmon qubit coherently couples to the strain of an HBAR acoustic mode via a disk of piezoelectric aluminum nitride patterned on the sapphire. Each qubit on the chip is nearly resonant with an HBAR mode of interest; in such a way, a pair of qubit-HBAR modes behaves like a single atom coupled to a phononic mode. The qubits were read out via on-chip microwave resonators using standard circuit quantum electrodynamics techniques Krantz et al. (2019); Reed et al. (2010). Our device is similar to those used in previous work generating mechanical Schrödinger cat states Bild et al. (2023) and for circuit quantum acoustodynamics Chu et al. (2017).
@@
where $\omega_{\rm r,q}$ are the readout and qubit frequency, $\hat{a}^{(\dagger)}$ is the photon annihilation (creation) operator, $\hat{\sigma}_{\rm z}$ is the qubit population operator, and $\chi$ is the qubit-state dependent frequency shift of the readout resonator. The coupling between the phonon and the qubit is described by a resonant Jaynes-Cummings interaction Chu et al. (2017), given by the Hamiltonian:
@@
The qubit’s state was measured by applying a weak probe tone on resonance with the readout resonator and monitoring the transmitted signal as a second tone was swept near the qubit frequency. The transmission spectrum at the readout frequency directly maps to the qubit occupation $\langle\hat{\sigma}_{\rm z}\rangle$; see Fig. 2(b). The asymmetry of the qubit spectrum is due to the finite photon population within the readout resonator Gambetta et al. (2006); Lachance-Quirion et al. (2017) and is well described by our theoretical model. Moreover, the narrow transparency window within the qubit spectrum results from the coherent swaps between the qubit and the phonon mode. In this work, the qubit and the phonon frequency were detuned by approximately $3.3$ MHz. See the discussion in the supplementary text for full details. However, as will be discussed below, coherent Rabi oscillations between the qubit and phonon modes do not provide a clear picture of the dynamics at high drive powers.
@@
Three distinct features can be observed within the two-tone spectrum when increasing the qubit drive power. First, the qubit linewidth is power-broadened Schuster et al. (2005); Gambetta et al. (2006). Large qubit drive powers increase the qubit decay rate due to increased stimulated emission. Less intuitive is the gradual disappearance and narrowing of the phonon-induced transparency window; see Fig. 2(c). The total linewidth of the transparency window is proportional to the phonon linewidth and demonstrates an inverse dependence on the drive power; see Fig. 2(d). This dependence is reminiscent of linewidth narrowing described by the Schawlow-Townes equation and is a known property of lasing states Gordon, Zeiger, and Townes (1955). These features can be understood by considering the schematic shown in Fig. 3(a).
As described before, the qubit and phonon modes coherently exchange excitations at the lowest qubit drive powers, which can be understood as a bosonic beam-splitter-type interaction. However, as the qubit drive power increases, the qubit is excited by the drive and coherently exchanges an excitation with the phonon mode. But before the phonon mode can exchange the excitation back to the qubit, the external drive re-excites the qubit. The qubit excitation can again be coherently swapped into the phonon mode, causing the phonon mode to climb the Fock state ladder. In the steady state, this process results in a phonon mode described by a large coherent state; i.e. the phonon mode starts lasing. A strong qubit drive allows the phonon mode to accept energy from the qubit but prevents the qubit from accepting energy from the phonon mode; see Fig. 3(a). This process also provides an intuitive understanding of the reduced visibility of the transparency window. As the drive power increases, the phonon mode no longer exchanges energy with the qubit; therefore, probing the qubit provides no information about the phonon mode.
@@
Figure 3: Phonon lasing observed through qubit ringdown dynamics. (A) Schematic of the coupled qubit-phonon system. The qubit acts as an artificial two-level atom, with ground state $|g\rangle$ and excited state $|e\rangle$. A coherent drive of strength $\Omega_{\rm q}$ drives the qubit between its ground and excited state. The qubit coherently interacts with the phonon mode at a rate $g_{\rm qb}$. The qubit is rapidly re-excited for strong pump powers, resulting in a one-directional energy flow into the phonon mode. This process results in phonon lasing. (B) Measured qubit ringdown for a qubit drive of -3.0 dBm. Blue line: the qubit drive is detuned from the HBAR by $250$ kHz with the phonon mode in the non-lasing state, decaying on a time scale of the $\sim 200$ ns decay of the qubit. Red line: The qubit drive is tuned directly on resonance with the HBAR mode, exciting it into the lasing state, exhibiting a dramatically longer, non-exponential decay due to re-excitation from the coherently excited phonon mode. The master equation simulation (dashed black line) is plotted over the data.
Finally, at the highest qubit drive powers, the qubit linewidth has been increased such that the rapid decay of the qubit population results in a self-quenching of phonon lasing Mu and Savage (1992). Counterintuitively this results in an upper lasing threshold above which the phonon mode is no longer lasing. Instead, the qubit incoherently exchanges excitations with the phonon mode, reducing the phonon amplitude. It should also be noted that the phonon mode statistics are no longer described as a coherent state above the upper threshold. This process has been described previously in the context of single-atom photon lasers McKeever et al. (2003) and is captured by our theoretical description; see the supplementary text.
@@
Figure 4: Unique signatures of single-atom lasing. (A) Measured gated qubit ringdown for a qubit drive of -12.0, -8.0, -6.0 and -2.0 dBm. With increasing drive power, the gated qubit ringdowns increase in amplitude and duration as the phonon mode population increases. (B) Measured gated qubit ringdown for a qubit drive of -2.0, 2.0, 4.0, and 8.0 dBm. For all measurements, the qubit drive is tuned directly on resonance with the HBAR mode. With increasing drive power, the gated qubit ringdowns decrease in amplitude and duration as the phonon mode population decreases above the self-quenching threshold. For all measurements in (A) and (B), the qubit drive is tuned directly on resonance with the HBAR mode. (C) Simulated phonon state population $\langle\hat{b}^{\dagger}\hat{b}\rangle$ as a function of qubit drive power. The phonon population is maximum at a drive power of approximately -2.0 dBm and decreases for drive powers above the upper lasing threshold. The individual trace colors in (A) and (B) correspond to the matching colors within the phonon population plot.
@@
Directly probing the phonon mode is not possible in the current experimental configuration since the readout is performed via the two-level system and not through a propagating photon mode Kuang et al. (2023); Behrle et al. (2023). Direct measurements of the Rabi oscillations between the qubit and phonon state could be possible and have been previously used to measure Fock-states in $\hbar$BAR devices Chu et al. (2018). However, this data would provide little clarity due to the short lifetime of our qubit and the multiplicity of simultaneous Rabi oscillation frequencies given by $2g_{\rm N}=2\sqrt{N}g_{\rm qb}$, which scale with the phonon mode amplitude $N$.Chu et al. (2018) Instead, we rely on the mismatch between the decay rate of the phonon mode $\tau\sim 25$ $\mu$s and that of the qubit. Using gated two-tone ringdown measurements, we can distinguish pure qubit decay from qubit decay driven by a highly excited phonon state. If the phonon mode is highly excited—in the absence of an external drive—the coherent Jaynes-Cummings interaction will continually drive the qubit, resulting in an extended relaxation of the qubit population compared to its intrinsic relaxation rate.
Gated two-tone measurements were performed, driving the qubit until the system reached its steady state; at this point, the drive was switched off using an RF switch. During the entire sequence, the frequency of the readout resonator is monitored using a vector network analyzer, averaging multiple traces triggered synchronously with the RF switch; see the supplementary text for more information. This measures the expectation value of the qubit population $\langle\hat{\sigma}_{\rm z}\rangle$ as a function of time, with a temporal resolution of $50$ ns.

### v2

@@ A Superconducting Single-Atom Phonon Laser
1Kavli Institute of Nanoscience, Delft University of Technology, PO Box 5046, 2600 GA Delft, The Netherlands
2ISIS (UMR 7006), Université de Strasbourg, 67000 Strasbourg, France
3Institute for Theory of Condensed Matter, Karlsruhe Institute of Technology, 76131 Karlsruhe, Germany
4Institute for Quantum Materials and Technology, Karlsruhe Institute of Technology, 76344 Eggenstein-Leopoldshafen, Germany
Email: clinton.potts@nbi.ku.dk
Email: g.a.steele@tudelft.nl
The development of quantum acoustics has enabled the cooling of mechanical objects to their quantum ground state, generation of mechanical Fock-states, and Schrödinger cat states. Such demonstrations have made mechanical resonators attractive candidates for quantum information processing, metrology, and macroscopic tests of quantum mechanics. Here, we experimentally demonstrate a direct quantum-acoustic equivalent of a single-atom laser. A single superconducting qubit coupled to a high-overtone bulk acoustic resonator is used to drive the onset of phonon lasing. We observe the absence of a sharp lower lasing threshold and characteristic upper lasing threshold, unique predictions of single-atom lasing. Lasing of an object with a $25$ $\mu$g mass represents a new regime of laser physics. It provides a possible tool for generating large amplitude coherent states in circuit quantum acoustodynamics, which is important for fundamental and quantum information applications.
@@ INTRODUCTION
The invention of the laser in the early 1960s has proven to be one of the most important breakthroughs in modern physics. The laser found its niche in the advancement of atomic physics, resulting in numerous groundbreaking discoveries such as spectroscopic measurements of atomic species Schawlow (1982, 1978), laser cooling of atoms Phillips and Metcalf (1982); Phillips (1998), and the realization of Bose-Einstein condensation Anderson et al. (1995); Davis et al. (1995). Beyond this, lasers are indispensable tools in areas like gravitational wave detection Abbott et al. (2016), biology Chaurasiya et al. (2016), chemistry McCreery (2005), and even in medicine Khalkhal et al. (2019). The advancements enabled by optical lasers have motivated the development of lasing states of mechanical vibrations — phonon lasersMahboob et al. (2013); Vahala et al. (2009); Pettit et al. (2019); Kuang et al. (2023); Behrle et al. (2023). Recent examples of phonon lasing have been realized with the center of mass motion of levitated microspheres or ions Vahala et al. (2009); Pettit et al. (2019); Kuang et al. (2023); Behrle et al. (2023), limiting the integrability of these coherent phonon sources. In other examples of phonon lasers in literature, the lasing state is the result of dispersive or dissipative optomechanical coupling Shandilya et al. (2021); Cui et al. (2021); Zhang et al. (2018); Grudinin et al. (2010); Potts et al. (2021); Parsa et al. (2024) which are well described as mechanical parametric instabilities Kippenberg and Vahala (2008) or via electron transport driven Kent et al. (2006) and optically driven semiconductor heterostructures Bron and Grill (1978); Wallentowitz et al. (1996); Camps et al. (2001); Tilstra, Arts, and De Wijn (2007).
The prospect of controlling and manipulating phonons at the quantum level has sparked the development of the field of circuit quantum acoustodynamics (cQAD), in which single superconducting qubits are coupled to bulk acoustic phonon modes Chu et al. (2017). Circuit QAD has seen rapid success with demonstrations such as the generation of arbitrary Fock states Chu et al. (2018), and Schrödinger cat states Bild et al. (2023), and the prospect of developing novel hybrid quantum systems Chu and Gröblacher (2020); Välimaa et al. (2022); Hann et al. (2019). However, as the acoustic mode is coupled only via the qubit in current cQAD platforms, it can be challenging to generate the large coherent phonon amplitudes required for creating large cat states and for phonon mode tomography Bild et al. (2023).
@@
Figure 1: Schematic of the on-chip $\hbar$BAR device. Rendering of the $\hbar$BAR device. The $\hbar$BAR device comprises two chips bounded in a flip-chip orientation. The top chip is 650 $\mu$m of sapphire, hosts the HBAR modes (red), and is coupled to the superconducting antenna using an aluminum nitride pad (red). The pocket-style transmon qubit (silver) is fabricated from NbTiN on the bottom silicon chip and coupled to the feedline via an on-chip readout resonator (blue).
@@
Figure 2: Qubit-induced acoustic linewidth narrowing. a Definition of modes and the coupling rates between the modes. The readout resonator $\hat{a}$ has input-output modes labelled $1$ and $2$. b Measured two-tone spectroscopy for drive power of -12.0 dBm set at room temperature. c Measured phonon-induced transparency window as a function of qubit drive power. Starting from the lowest curve, drive powers are -12.0, -6.0, and -2.0 dBm set at room temperature, respectively. With increasing power, two features can be noticed. The transparency window reduces in depth, and the full-width half-maximum linewidth narrows. d Extracted experimental full-width half-maximum of the transparency window as a function of qubit drive power. Colored points match the corresponding curves in c.
@@
In this article, we demonstrate the generation of large coherent phonon populations in cQAD using a direct phononic analog of a single-atom optical laser McKeever et al. (2003); Dubin et al. (2010); An et al. (1994). A single two-level artificial atom couples to a bulk on-chip phonon mode, generating a large amplitude phonon coherent state through the process of single-atom lasing. Similar to conventional lasers, in single-atom lasers, the excitations in the lasing mode are dominantly produced by stimulated emission Mu and Savage (1992), enabling a potential route to large coherent state preparation in cQAD. The lasing is observed through a dramatically extended ringdown of the artificial atom excited by the phonon laser and demonstrates the nearly threshold-free nature of single-atom lasers. We also observe a unique predicted feature of single-atom lasers, an upper lasing threshold Dubin et al. (2010); Mu and Savage (1992); Ashhab et al. (2009). Our experimental observations agree with the theoretical and numerical models we use to characterize the system.
Finally, the phonon mode’s bulk nature may allow the integration of color centers or quantum dots, enabling strain engineering or coherent mechanical driving for future hybrid quantum systems. Moreover, the coherent state generated via phonon lasing may be used as an efficient displacement pulse for generating large Schrödinger cat states Bild et al. (2023) or squeezed states Marti et al. (2024). Generating large mechanical cat states has exciting potential applications for future macroscopic tests of quantum mechanics Schrinski et al. (2023).
@@
Our device comprises a flip-chip $\hbar$BAR architecture with two bonded device chips Franse et al. (2024); Bild et al. (2023); von Lüpke et al. (2024, 2022); Chu et al. (2018, 2017). We have implemented a fully on-chip integration, which can multiplex different devices on a single silicon chip Välimaa et al. (2022); Crump, Välimaa, and Sillanpää (2023). A single feedline is coupled to the individual superconducting fixed-frequency transmon qubits. The transmon qubits were fabricated from niobium-titanium nitride for the bulk structures with aluminum Josephson junctions. A sapphire substrate, $650$ $\mu$m thick, was positioned above each transmon qubit and bonded to the silicon substrate, see Fig. 1. The sapphire chip supports a set of longitudinal high-overtone bulk acoustic wave resonances (HBARs) separated by a free spectral range of $8.54$ MHz. The electric field of a transmon qubit coherently couples to the strain of an HBAR acoustic mode via a disk of piezoelectric aluminum nitride patterned on the sapphire. Each qubit on the chip is nearly resonant with an HBAR mode of interest; in such a way, a pair of qubit-HBAR modes behaves like a single atom coupled to a phononic mode. The qubits were read out via on-chip microwave resonators using standard circuit quantum electrodynamics techniques Krantz et al. (2019); Reed et al. (2010). Our device is similar to those used in previous work generating mechanical Schrödinger cat states Bild et al. (2023) and for circuit quantum acoustodynamics Chu et al. (2017).
@@
where $\omega_{\rm r,q}$ are the readout and qubit frequency, $\hat{a}^{(\dagger)}$ is the photon annihilation (creation) operator, $\hat{\sigma}_{\rm z}$ is the qubit Pauli-z operator, and $\chi$ is the qubit-state dependent frequency shift of the readout resonator. The coupling between the phonon and the qubit is described by a resonant Jaynes-Cummings interaction Chu et al. (2017), given by the Hamiltonian:
@@
The qubit’s state was measured by applying a weak probe tone on resonance with the readout resonator and monitoring the transmitted signal as a second tone was swept near the qubit frequency. The transmission spectrum at the readout frequency directly maps to the qubit occupation $\langle\hat{\sigma}_{+}\hat{\sigma}_{-}\rangle$; see Fig. 2(b). The asymmetry of the qubit spectrum is due to the finite photon population within the readout resonator Gambetta et al. (2006); Lachance-Quirion et al. (2017) and is well described by our theoretical model. Moreover, the narrow transparency window within the qubit spectrum results from the weak hybridization between the qubit and the phonon mode. In this work, the qubit and the phonon frequency were detuned by approximately $3.3$ MHz. See the discussion in the supplementary text for full details. However, as will be discussed below, coherent energy exchange between the qubit and phonon modes do not provide a clear picture of the dynamics at high drive powers.
@@
Three distinct features can be observed within the two-tone spectrum when increasing the qubit drive power. First, the qubit linewidth is power-broadened Schuster et al. (2005); Gambetta et al. (2006). Large qubit drive powers increase the qubit decay rate due to increased stimulated emission. Less intuitive is the gradual disappearance and narrowing of the phonon-induced transparency window; see Fig. 2(c). The total linewidth of the transparency window is proportional to the phonon linewidth and demonstrates an inverse dependence on the drive power; see Fig. 2(d). This dependence is reminiscent of linewidth narrowing described by the Schawlow-Townes equation and is a known property of lasing states Gordon, Zeiger, and Townes (1955). However, it is important to note that, here, we observe a narrowing of the induced transparency window and not a narrowing of the emission spectrum. These features can be understood by considering the schematic shown in Fig. 3(a).
The build-up of the phonon population can be understood by considering the different processes through which excitations can be transferred between the phonon and the qubit. At low driving powers, the linewidth broadening of the qubit due to the microwave drive is small, and the phonon mode and qubit hybridize weakly. The weak hybridization combined with the microwave drive initiates the building up of the phonon population. A balance between the rate at which the qubit is excited and de-excited, the rate at which excitations are transferred between qubit and phonon mode and the stochastic phonon losses yield a steady-state occupation of the phonon mode. For low drive powers and thus slow qubit excitation rates, phonon loss dominates, preventing the build-up of the phonon population. As the qubit excitation rate increases, transitions between states of the Jaynes-Cummings ladder with higher phonon numbers can be efficiently driven. Since the rate at which excitations are transferred between states of the Jaynes-Cummings ladder depends on the phonon number of the involved states, as the average population of the phonon mode increases, the rate at which excitations are transferred correspondingly increases; this is the origin of acoustic stimulated emission Hofheinz et al. (2008). In the steady state, this process, balanced by stochastic phonon losses, results in a phonon mode described by a large coherent state, i.e. the phonon mode starts lasing. Such processes also provide an intuitive understanding of the reduced visibility of the transparency window. As the drive power increases, the linewidth of the qubit becomes increasingly broad, and the coupling between qubit and phonon is not strong enough to overcome the qubit loss; therefore, the qubit spectroscopy provides limited information about the phonon mode.
@@
Figure 3: Phonon lasing observed through qubit ringdown dynamics. a Schematic of the energy levels for the weakly hybridized phonon-qubit system. The qubit acts as an artificial two-level atom, with ground state $|g\rangle$ and excited state $|e\rangle$. A coherent drive of strength $\Omega_{\rm q}$ drives the qubit between its ground and excited state. The rate at which excitation transfer processes $|e,N-1\rangle\rightarrow|g,N\rangle$ occur scales as $\sqrt{N}$, where $N$ is the total number of excitations, while the phonon relaxation scales linearly with $N$. The qubit is rapidly re-excited for strong pump powers, resulting in a build-up of phonon excitations. This process results in phonon lasing. b Measured qubit ringdown for a qubit drive of -3.0 dBm. Blue line: the qubit drive is detuned from the HBAR by $250$ kHz with the phonon mode in the non-lasing state, decaying on a time scale of the $\sim 200$ ns decay of the qubit. Red line: The qubit drive is tuned directly on resonance with the HBAR mode, exciting it into the lasing state, exhibiting a dramatically longer, non-exponential decay due to re-excitation from the coherently excited phonon mode. The master equation simulation (dashed black line) is plotted over the data.
@@
Figure 4: Unique signatures of single-atom lasing. a Measured gated qubit ringdown for a qubit drive of -12.0, -8.0, -6.0, and -2.0 dBm. With increasing drive power, the gated qubit ringdowns increase in amplitude and duration as the phonon mode population increases. b Measured gated qubit ringdown for a qubit drive of -2.0, 2.0, 4.0, and 8.0 dBm. The qubit drive is tuned directly on resonance with the HBAR mode for all measurements. With increasing drive power, the gated qubit ringdowns decrease in amplitude and duration as the phonon mode population decreases above the self-quenching threshold. For all measurements in a and b, the qubit drive is tuned directly on resonance with the HBAR mode. c Simulated phonon state population $\langle\hat{b}^{\dagger}\hat{b}\rangle$ as a function of qubit drive power. The phonon population is at its maximum at a drive power of approximately -2.0 dBm and decreases for drive powers above the upper lasing threshold. The colored data points indicate the corresponding trace colors in a and b.
At the highest qubit drive powers, the qubit linewidth has been increased such that the rapid decay of the qubit results in a self-quenching of phonon lasing Mu and Savage (1992). This results in an upper lasing threshold above which the phonon mode is no longer lasing. Above this threshold, the qubit undergoes rapid Rabi oscillations due to the strong microwave drive and, therefore, cannot exchange excitations with the phonon mode, reducing the phonon amplitude. It should also be noted that the phonon mode statistics are no longer described as a coherent state above the upper threshold. This process has been described previously in the context of single-atom photon lasers McKeever et al. (2003) and is captured by our theoretical description; see the supplementary text.
@@
Directly probing the phonon mode is not possible in the current experimental configuration since the readout is performed via the two-level system and not through a propagating photon mode Kuang et al. (2023); Behrle et al. (2023). Direct measurements of the Rabi oscillations between the qubit and phonon state have been previously used to measure Fock-states in $\hbar$BAR devices Chu et al. (2018). However, this data would not be possible due to the short lifetime of our qubit and the multiplicity of simultaneous Rabi oscillation frequencies between two Fock states is given by $2g_{\rm N}=2\sqrt{N}g_{\rm qb}$, which scale with the phonon Fock number $N$Chu et al. (2018). Instead, we rely on the mismatch between the decay rate of the phonon mode $\tau\sim 25$ $\mu$s and that of the qubit. Using gated two-tone ringdown measurements, we can distinguish pure qubit decay from qubit decay driven by a highly excited phonon state. If the phonon mode is highly excited—in the absence of an external drive—the coherent Jaynes-Cummings interaction will continually drive the qubit, resulting in an extended relaxation of the qubit population compared to its intrinsic relaxation rate.
Gated two-tone measurements were performed, driving the qubit until the system reached its steady state; at this point, the drive was switched off using an RF switch. During the entire sequence, the frequency of the readout resonator is monitored using a vector network analyzer, averaging multiple traces triggered synchronously with the RF switch; see the supplementary text for more information. This measures the expectation value of the qubit population $\langle\hat{\sigma}_{+}\hat{\sigma}_{-}\rangle$ as a function of time, with a temporal resolution of $50$ ns.